# Divided We Fall: Defending Against Adversarial Attacks via Soft-Gated Fractional Mixture-of-Experts with Randomized Adversarial Training

### v2

@@ Abstract
Machine learning is a powerful tool enabling full automation of a huge number of tasks without explicit programming. Despite recent progress of machine learning in different domains, these models have shown vulnerabilities when they are exposed to adversarial threats. Adversarial threats aim to hinder the machine learning models from satisfying their objectives. They can create adversarial perturbations, which are imperceptible to humans’ eyes but have the ability to cause misclassification during inference. In this paper, we propose a defense system, which devises an adversarial training module within mixture-of-experts architecture to enhance its robustness against white-box evasion attacks. In our proposed defense system, we use nine pre-trained classifiers (experts) with ResNet-18 as their backbone. During end-to-end training, the parameters of all experts and the gating mechanism are jointly updated allowing further optimization of the experts. Our proposed defense system outperforms state-of-the-art MoE-based defenses under strong white-box FGSM and PGD evaluation on CIFAR-10 and SVHN.
@@
organization=Computer Science, University of New Brunswick, city=Fredericton, postcode=E3B 5A3, country=Canada
@@ 1 Introduction
Machine learning (ML) is a cornerstone of artificial intelligence (AI), which has been widely used in numerous applications. In computer vision, it enables face recognition and video surveillance [26, 27]. In health care, it facilitates disease diagnosis, drug repurposing, and personalized treatments [17, 44]. In natural language processing (NLP), ML enhances the quality of language translation [31, 39]. ML also empowers Large Language Models (LLMs) for code-related tasks such as code generation and understanding [22]. In cybersecurity, ML eases threat detection and manages system’s vulnerabilities [3, 56, 9]. Moreover, ML also plays an important role in other domains such as finance, speech processing, recommendation systems, autonomous vehicles, among others. However, those wide range of applications of machine learning has introduced opportunities for adversaries to exploit natural vulnerabilities of these models and launch adversarial attacks to threaten the integrity, privacy, and availability of these models [41, 8, 29, 25].
Adversarial machine learning (AML) is at intersection of cybersecurity and machine learning and aims to investigate attacks on machine learning models, and design defense systems to counter these attacks. The adversary may have some knowledge about the victim’s model and/or its training data, which facilitates the adversarial sample generation [41, 8, 29]. Figure 1 shows how an adversarial sample looks like on an MNIST [32] sample when the adversary has a perfect knowledge about the model’s parameters. Samples like this can cause model’s misclassification with a high confidence during test time.
@@
The rivalry between governments and industrial corporations has given rise to adversarial attacks causing global concerns about adversarial machine learning. For instance, the government of United States of America has established Bipartisan AI task force, which are responsible for appropriate use of AI. They introduced AI guidelines such as accuracy, reliability, robustness, privacy, transparency, and safety. They also labeled generative and adversarial AI as a threat to privacy and confidentiality [7]. Moreover, the Department of Homeland Security (DHS) has considered evasion attacks and generative AI as a national threat violating security and privacy of individuals [30]. These insights highlight the need for secure AI systems and robust ML models.
Trustworthy AI is an important domain of AI ensuring that AI models remain lawful, ethical, robust, and efficient. Robust machine learning is an important branch in trustworthy AI which aims to make the machine learning models robust against adversarial threats while remain functional in non-adversarial environments [25]. Moreover, the main focus of Robust ML is to develop defense systems that detect and/or mitigate adversarial threats while preserving their efficiency and performance to a certain extent. These defense systems could either reside in the models or serve as an external entity protecting ML models [41, 8, 29, 25].
Although adversarial machine learning encompasses multiple threat categories, including poisoning, model extraction, and privacy attacks, this study focuses specifically on inference-time robustness against white-box evasion attacks. Our evaluation therefore targets gradient-based adversarial examples generated with full model access. Training-time attacks and broader privacy or extraction threats are outside the scope of this work and are left for future investigation.
@@
  • Our proposed defense mechanism devises an adversarial training module within MoE, which outperforms related state-of-the-art defense systems in terms of both clean accuracy and robustness.
  • Our model maintains high accuracy, while significantly outperforms undefended models in terms of robustness.
@@
In Section 2, we illustrate the background problem setting. In Section 3, we introduce similar defense systems. In Section 4, we present the methodology and workflow of our defense. In Section 5, we demonstrate our experimental results. In Section 6, we investigate our proposed defense system limitations, potential future directions, and finally conclude our work.
@@ 2.1 Mixture-of-Experts
MoE is an ensemble architecture using multiple experts, a gating mechanism to assign weights, and a router to aggregate results [37, 10, 43]. MoEs can be categorized based on their gating mechanism, routing topology, training strategy, and experts activation, which is shown in Figure 2.
@@
Based on gating mechanism, MoE could be either soft or hard; in soft MoE the gating assigns non-zero weights to each expert forming fractional activation of experts, while in hard MoE, the gating mechanism chooses a subset of experts and assign wights to them making experts activation as sparse, or single if only one of them is chosen [37, 10, 43].
Training strategy could be either joint or frozen. In the first one, both router and experts are trained simultaneously, while the latter one uses pretrained experts and freezes their weights during training of the router [37, 10, 43].
Based on the routing topology we have flat and hierarchical routing schemes. In the flat MoE, all the experts reside in a single layer and the router chooses from them. In the hierarchical topology, the first router groups the experts into subgroups, and, then, the second one chooses some experts from these subgroups [37, 10, 6].
@@ 2.2.1 Attacker’s Knowledge
Attacker’s knowledge shows the access level of the adversary to model’s parameters, gradients, architecture, and training data. This knowledge is categorized into white-box, black-box, and gray-box based on the adversary’s access level [41, 8].
@@
Adversaries with white-box knowledge are the most effective and destructive attackers. These attackers have a full access to model’s architecture, parameters, and gradients during training with full knowledge on the training data; this enables the attackers to generate adversarial samples based on the gradients and training data, which makes them very powerful. However, in a real-world scenario, adversaries can’t achieve such an access, making them more suitable for research and testing the models’ robustness [41, 8].
@@
In black-box scenario, unlike white-box, the adversaries have no knowledge about the model and the presence of any defense system. The adversaries can only give inputs and receive the corresponding outputs [41, 8, 33].
@@
Gray box knowledge lies between white box knowledge and black box knowledge. The adversary has some degree of knowledge about the target system, which is not as perfect as white box, and is not as scarce as black box knowledge [41, 8].
@@ 2.2.2 Attacker’s specificity
Based on specificity, each attack could either be targeted or untargeted. In targeted attacks, the adversary aims to degrade the model’s performance when it faces specific patterns/inputs, while in untargeted attacks, also named to as indiscriminate, the goal is to cause any misclassification [41, 8].
@@ 2.2.3 Attacker’s objective
Based on the objective, the adversary can follow degrade, breach, or both scenarios. In the first scenario, the adversary aims to harm the system’s performance and/or violets its integrity. In breach, the adversary aims to explore the model in order to steal sensitive information, without affecting the model’s performance [41, 8].
@@ 2.2.4 Attacker’s type
An evasion attack targets the model at test time without affecting the system’s integrity on clean data. Evasion attacks are the most practical attack types aiming to degrade the model’s performance by causing misclassification during testing [41, 8, 15]. For instance attacks such as Fast Gradient Sign Method (FGSM) [16], Projected Gradient Descent (PGD) [36], and Carlini & Wagner (C&W) [11] make use of the model’s gradient during training to generate samples which evade the classifier during the test time. Equations 3 and 4, in Section 5, show how an adversarial sample is generated in FGSM and PGD, respectively.
@@
Poisoning attacks compromise a model’s integrity by corrupting training data or embeddings, either through label manipulation or clean label attacks [41, 8, 55, 5, 59, 65]. Exploratory attacks aim to violate privacy of the system. These attacks probe the target model to gain information about model’s parameters, training data, and algorithms on without affecting the performance and integrity of the model [46, 62, 50]. Moreover, physical attacks are launched on the models, which are deployed on a special physical hardware such as FPGA, GPU, and edge devices to exploit vulnerabilities emerged due to the increased attack surface in the physical layer [45, 20]. In this research work, we only focus on defending against evasion attacks to assess the resilience of our proposed approach under severe white-box attacks.
Figure 3 shows an overview of adversarial threat model.
@@ 3 Related Works
Adversarial training is a defense strategy aiming to train the model on both benign and adversarial samples to increase generalization capability of the model enabling it to correctly predict both benign and adversarial inputs. However, this method often comes with a trade-off, where the model sacrifices it accuracy on the benign data to increase its robustness against adversarial inputs [4, 64, 63, 23].
@@
[42, 40] investigate the adversarial robustness of MoEs. In [18, 60], some defense systems are proposed, which aim to protect the architecture of MoEs. Defenses such as adversarial mixture-of-expert (ADVMoE) [61] and synergy-of-experts (SoE) [13] make use of the ensemble architecture of MoEs to defend against adversarial threats. Both ADVMoE and SoE use adversarial training techniques, while training their MoE to increase the robustness of their model.
Table 1 presents the differences between the category of MoE used in our defense system and other competitors.
Table 1: MoE type of DWF vs. other competitors - gating mechanism (GM), routing topology (RT), training strategy (TS), and experts activation (EA).
Defense GM RT TS EA
SoE Hard Flat Joint Single
ADVMoE Hard Flat Joint Sparse
DWF Soft Flat Pre-trained+Joint Fractional
@@ 4 Methodology
The gating function is implemented as a lightweight learnable router that takes the same input $x$ as the experts and produces a soft assignment vector using a small multilayer perceptron followed by a softmax normalization. The final prediction is obtained via a weighted sum of all expert logits (soft routing), rather than hard or sparse selection. All experts and the gating network are trained jointly in an end-to-end manner using the standard cross-entropy loss. No auxiliary routing objectives, load-balancing losses, temperature scaling, or sparsity regularization are used. This simple design avoids additional hyperparameters and improves reproducibility.

### v3

@@ Abstract
Machine learning is a powerful tool enabling full automation of a huge number of tasks without explicit programming. Despite recent progress of machine learning in different domains, these models have shown vulnerabilities when they are exposed to adversarial threats. Adversarial threats aim to hinder the machine learning models from satisfying their objectives. They can create adversarial perturbations, which are imperceptible to humans’ eyes but have the ability to cause misclassification during inference. In this paper, we propose a defense system, which devises an adversarial training module within mixture-of-experts architecture to enhance its robustness against white-box evasion attacks. In our proposed defense system, we use nine pre-trained classifiers (experts) with ResNet-18 as their backbone. During end-to-end training, the parameters of all experts and the gating mechanism are jointly updated allowing further optimization of the experts. Our proposed defense system outperforms prior MoE-based defenses under strong white-box FGSM and PGD evaluation on CIFAR-10 and SVHN. The use of multiple experts increases training time and compute relative to single-network baselines; however, inference scales approximately linearly with the number of experts and is substantially cheaper than training.
@@
organization=Trustworthy and Secure AI (TSAI) Lab, Faculty of Computer Science, University of New Brunswick, city=Fredericton, country=Canada
@@ 1 Introduction
Machine learning (ML) is a cornerstone of artificial intelligence and is widely deployed in applications such as computer vision [18], healthcare [31], natural language processing [26], cybersecurity [3], and autonomous systems. Despite its success, the widespread adoption of ML has exposed inherent vulnerabilities that adversaries can exploit through adversarial attacks, threatening the integrity, privacy, and availability of deployed systems [28]. Adversarial machine learning (AML) lies at the intersection of cybersecurity and ML, studying attacks that manipulate model behavior and developing defenses to mitigate them. Depending on the threat model, adversaries may possess partial or full knowledge of the model and training data, enabling the generation of adversarial examples that induce misclassification at inference time [7]. Figure 1 shows how an adversarial sample looks like on an MNIST [20] sample when the adversary has a perfect knowledge about the model’s parameters. Samples like this can cause model’s misclassification with a high confidence during test time.
@@
These risks have elevated adversarial ML to a global concern. Government and regulatory bodies have identified adversarial and generative AI as threats to privacy, security, and public safety, emphasizing robustness, reliability, and transparency as key requirements for trustworthy AI systems [6]. Such concerns underscore the urgent need for robust ML models that remain reliable under adversarial conditions. Robust machine learning, a core component of trustworthy AI, focuses on designing defense mechanisms that mitigate adversarial threats while preserving performance in benign settings [17]. Among emerging approaches, mixture-of-experts (MoE) architectures offer a promising direction by enabling specialization and adaptive routing, which can enhance resilience to adversarial perturbations while maintaining efficiency.
We consider a white-box adversarial setting, where the attacker has full access to the model parameters, gradients, architecture, and training procedure. The attacker’s objective is to induce misclassification at inference time by crafting adversarial examples under an $\ell_{\infty}$ norm. The attacker has white-box knowledge and optimizes perturbations accordingly. We do not consider data poisoning, model extraction, or privacy attacks in this work; our evaluation focuses exclusively on evasion robustness.
@@
Methodological contributions:
@@
Empirical findings:
  • Under a white-box $\ell_{\infty}$ threat model with FGSM and multi-step PGD attacks, DWF achieves higher robust accuracy than prior MoE-based defenses (ADVMoE, SoE) on CIFAR-10 and SVHN.
  • DWF preserves high clean accuracy (¿90%) while improving robustness relative to the defended baselines.
  • Robustness improvements remain consistent under stronger attacks, including multi-step and adaptive PGD, demonstrating stable performance across evaluation protocols.
In Section 2, we illustrate the background problem setting. In Section 3, we introduce similar defense systems. In Section 4, we present the methodology and workflow of our defense. In Section 5, we demonstrate our experimental results. In Section 6, we investigate our proposed defense system limitations and potential future directions. Finally, in Section 7, conclude our work.
@@ 2.1 Mixture-of-Experts
MoE is an ensemble architecture using multiple experts, a gating mechanism to assign weights, and a router to aggregate results [24]. MoEs can be categorized based on their gating mechanism, routing topology, training strategy, and experts activation, which is shown in Figure 2.
@@
Based on gating mechanism, MoE could be either soft or hard; in soft MoE the gating assigns non-zero weights to each expert forming fractional activation of experts, while in hard MoE, the gating mechanism chooses a subset of experts and assign wights to them making experts activation as sparse, or single if only one of them is chosen [30].
Training strategy could be either joint or frozen. In the first one, both router and experts are trained simultaneously, while the latter one uses pretrained experts and freezes their weights during training of the router [24].
Based on the routing topology we have flat and hierarchical routing schemes. In the flat MoE, all the experts reside in a single layer and the router chooses from them. In the hierarchical topology, the first router groups the experts into subgroups, and, then, the second one chooses some experts from these subgroups [24].
@@
Our design adopts soft routing with fractional expert activation in a flat MoE architecture to balance robustness, stability, and computational cost. Soft gating enables smooth expert weighting, avoiding brittle decision boundaries associated with hard routing and improving gradient flow during adversarial training. Fractional activation allows multiple experts to contribute to each prediction, increasing functional diversity and reducing single-expert failure under adversarial perturbations, at the cost of moderately increased computation compared to sparse activation. We choose a flat architecture to maintain simplicity and reproducibility, while still enabling specialization through independent experts. Overall, these choices prioritize robustness and graceful degradation under attack, while keeping the computational overhead manageable.
@@ 2.2.1 Attacker’s Knowledge
Attacker’s knowledge shows the access level of the adversary to model’s parameters, gradients, architecture, and training data. This knowledge is categorized into white-box, black-box, and gray-box based on the adversary’s access level [28].
@@
Adversaries with white-box knowledge are the most effective and destructive attackers. These attackers have a full access to model’s architecture, parameters, and gradients during training with full knowledge on the training data; this enables the attackers to generate adversarial samples based on the gradients and training data, which makes them very powerful. However, in a real-world scenario, adversaries can’t achieve such an access, making them more suitable for research and testing the models’ robustness [28].
@@
In black-box scenario, unlike white-box, the adversaries have no knowledge about the model and the presence of any defense system. The adversaries can only give inputs and receive the corresponding outputs [21].
@@
Gray box knowledge lies between white box knowledge and black box knowledge. The adversary has some degree of knowledge about the target system, which is not as perfect as white box, and is not as scarce as black box knowledge [28].
@@ 2.2.2 Attacker’s specificity
Based on specificity, each attack could either be targeted or untargeted. In targeted attacks, the adversary aims to degrade the model’s performance when it faces specific patterns/inputs, while in untargeted attacks, also named to as indiscriminate, the goal is to cause any misclassification [7].
@@ 2.2.3 Attacker’s objective
Based on the objective, the adversary can follow degrade, breach, or both scenarios. In the first scenario, the adversary aims to harm the system’s performance and/or violets its integrity. In breach, the adversary aims to explore the model in order to steal sensitive information, without affecting the model’s performance [28].
@@ 2.2.4 Attacker’s type
An evasion attack targets the model at test time without affecting the system’s integrity on clean data. Evasion attacks are the most practical attack types aiming to degrade the model’s performance by causing misclassification during testing [28]. For instance attacks such as Fast Gradient Sign Method (FGSM) [12], Projected Gradient Descent (PGD) [23], and Carlini & Wagner (C&W) [8] make use of the model’s gradient during training to generate samples which evade the classifier during the test time. Equations 3 and 4, in Section 5, show how an adversarial sample is generated in FGSM and PGD, respectively.
@@
Poisoning attacks compromise a model’s integrity by corrupting training data or embeddings, either through label manipulation or clean label attacks [40]. Exploratory attacks aim to violate privacy of the system. These attacks probe the target model to gain information about model’s parameters, training data, and algorithms on without affecting the performance and integrity of the model [45]. Moreover, physical attacks are launched on the models, which are deployed on a special physical hardware such as FPGA, GPU, and edge devices to exploit vulnerabilities emerged due to the increased attack surface in the physical layer [32]. In this research work, we only focus on defending against evasion attacks to assess the resilience of our proposed approach under severe white-box attacks. Figure 3 shows an overview of adversarial threat model.
@@
We evaluate inference-time robustness under a white-box $\ell_{\infty}$ threat model, where the adversary constructs an adversarial example $x_{\text{adv}}=x+\delta$ subject to $\|\delta\|_{\infty}\leq\epsilon$, with the goal of inducing misclassification. In our experiments, we report results under FGSM (single-step) and PGD (multi-step) attacks (including Multi-Step PGD and Adaptive PGD) using the settings summarized in Table 4. After each update step, the perturbation is projected back into the $\ell_{\infty}$ ball around the clean input, $x-\epsilon\leq x_{\text{adv}}\leq x+\epsilon$, and valid pixel bounds are enforced by clipping the perturbed sample to the input range $[0,1]$ using $x_{\text{adv}}=\texttt{clamp}(x_{\text{adv}},0,1)$. In this work, ”imperceptible” perturbations are defined operationally by the chosen $\ell_{\infty}$ budgets ($\epsilon\in\{0.01,0.02,\ldots,0.09\}$ for FGSM and $\epsilon=\frac{8}{255}$ for PGD-based attacks), which correspond to standard small-magnitude perturbations used in robustness evaluations.
@@ 3 Related Works
Adversarial training is a defense strategy aiming to train the model on both benign and adversarial samples to increase generalization capability of the model enabling it to correctly predict both benign and adversarial inputs [4]. However, this method often comes with a trade-off [46], where the model sacrifices it accuracy on the benign data to increase its robustness against adversarial inputs [47].
@@
[29] and [27] investigate the adversarial robustness of MoEs. In [13] and [43], some defense systems are proposed, which aim to protect the architecture of MoEs. Defenses such as adversarial mixture-of-expert (ADVMoE) [44] and synergy-of-experts (SoE) [10] make use of the ensemble architecture of MoEs to defend against adversarial threats. Both ADVMoE and SoE use adversarial training techniques, while training their MoE to increase the robustness of their model.
DWF is not a direct recombination of MoE and standard adversarial training as used in prior MoE defenses (e.g., SoE, ADVMoE). Concretely, DWF differs along the three axes the MoE pipeline exposes: training pipeline, routing, and activation. Unlike SoE/ADVMoE, which use hard routing with single/sparse expert activation under standard adversarial training, DWF adopts soft gating with fractional expert activation (weighted combination of all experts) and trains the system using a pretrain-then-joint scheme with randomized adversarial training during training. Table 1 presents the differences between our proposed approach and other competitors.
Table 1: MoE type of DWF vs. other competitors - gating mechanism (GM), routing topology (RT), training strategy (TS), experts activation (EA), and adversarial training (AT).
Defense GM RT TS EA AT type
SoE Hard Flat Joint Single Standard
ADVMoE Hard Flat Joint Sparse Standard
DWF Soft Flat Pre-trained+Joint Fractional Randomized
@@
While mixture-of-experts defenses are the primary architectural focus of this work, it is important to position DWF relative to widely used non-MoE robustness baselines. Standard adversarial training methods improve robustness by optimizing a single model against worst-case perturbations. These approaches represent strong and widely adopted robustness baselines, but they lack explicit expert specialization and input-dependent routing.
Our evaluation includes adversarially trained ResNet-18 models (FGSM-AT and PGD-AT) as representative non-MoE baselines. The results show that DWF achieves higher robust accuracy while preserving stronger clean accuracy compared to these single-network defenses. The primary comparison with MoE-based methods (SoE and ADVMoE) is intentional, as it enables controlled architectural comparisons that isolate the impact of soft gating, fractional expert activation, and randomized adversarial training within the MoE paradigm.
To ensure that the observed robustness does not stem from gradient masking or obfuscated gradients, all attacks are implemented in a true white-box setting and differentiate through the full mixture-of-experts architecture, including all experts, the gating network, and the final loss. Perturbations are optimized end-to-end with respect to the complete model, preventing gradient obfuscation through routing or expert specialization. Robust accuracy decreases monotonically as attack strength increases (larger $\epsilon$ and more PGD iterations), which is inconsistent with gradient masking behavior. Moreover, inference in DWF is deterministic; therefore, expectation over transformation (EoT) with multiple samples yields identical gradients and attack outcomes, confirming the absence of stochastic masking.
@@ 4 Methodology
During adversarial training, adversarial examples are generated with respect to the full end-to-end loss of the mixture-of-experts model, allowing gradients to propagate through all experts and the gating mechanism. No random restarts are used during training. For FGSM, a single-step gradient update is applied using a randomly sampled perturbation magnitude. For PGD, iterative updates are performed with projection onto the $\ell_{\infty}$ ball around the clean input after each step. In all cases, perturbed samples are clipped to the valid input range $[0,1]$ to ensure valid pixel values.
@@
The gating function is implemented as a lightweight learnable router that takes the same input $x$ as the experts and produces a soft assignment vector using a small multilayer perceptron followed by a softmax normalization. The router first flattens the input image and then applies four fully connected layers with sizes 1024, 512, 128, and the number of experts, respectively. ReLU activation is used after each hidden layer, and the final Softmax layer produces non-negative routing weights that sum to one. The router is trained jointly with the experts and does not share parameters with the expert backbones.
The final prediction is obtained via a weighted sum of all expert logits (soft routing), rather than hard or sparse selection. All experts and the gating network are trained jointly in an end-to-end manner using the standard cross-entropy loss. No auxiliary routing objectives, load-balancing losses, temperature scaling, or sparsity regularization are used. This simple design avoids additional hyperparameters and improves reproducibility.
@@
Randomization is applied only during training, where adversarial parameters are sampled from predefined ranges (e.g., $\epsilon$ for FGSM and iteration counts for PGD) to expose the model to diverse perturbation strengths. This improves robustness generalization. During inference, no randomness is used: the gating mechanism and experts operate deterministically, and the model produces a fixed output for a given input. Therefore, expectation over transformation (EoT) is not required at test time, and using EoT samples >1>1 yields the same gradients and attack outcomes; thus we set EoT samples =1=1 in the adaptive evaluation.
Mixture-of-experts models may suffer from expert collapse, where routing concentrates on a small subset of experts. Our design mitigates this risk through architectural choices rather than explicit regularization. We employ soft gating with fractional activation, which assigns a normalized weight to every expert for each input, ensuring that predictions are formed through a weighted combination rather than exclusive selection. This contrasts with hard or sparse routing, where only a few experts are active and collapse is more likely. Furthermore, experts are pretrained on different input regimes (benign, FGSM, and PGD), encouraging functional diversity and reducing the likelihood that a single expert dominates across all inputs. While we do not incorporate load-balancing losses or diversity regularizers, these design choices provide a theoretical basis for promoting balanced expert utilization.
Fractional activation evaluates all experts using continuous routing weights, resulting in a theoretical increase in inference cost that scales approximately linearly with the number of experts relative to a single backbone of similar capacity. This design avoids discrete routing decisions and preserves smooth gradient flow, which is beneficial in adversarial settings. While this may increase computational requirements compared to sparse expert selection, it reflects a deliberate trade-off favoring robustness and stability over minimal inference cost.